# Competing color superconductivity and color Kondo effect in quark matter

### v1

@@ IV Summary of Phases
We proceed to discuss the various phases system possesses, as functions of the RG invariant parameter $d$. As shown in Fig. 2, for all values of $d\in\mathbb{R}$ and for $0<\delta<\frac{1}{2}$ the impurity quark is completely screened by a multiparticle cloud of bulk superconducting quarks. In this regime, the Kondo fluctuations overwhelm the bulk superconducting order and in additions to the bulk mass gap $\Delta$ a new scale $T_{K}$, the Kondo Temperature, is dynamically generated which characterizes the screening of the impurity quark and the local impurity density of states which attains the familiar Lorentzian form. As $\delta$ increases past $\frac{1}{2}$, superconducting pairing increases and competes with Kondo screening, causing a bound state to emerge below the bulk mass gap. In the range $\frac{1}{2}<\delta<\frac{5}{4}$, the bound‐state energy is negative, and the impurity is screened by a single‐particle mode. At $\delta=\frac{5}{4}$, the bound level crosses zero, and for $\frac{5}{4}<\delta<2$, its positive energy makes the unscreened impurity lower in energy, marking a first-order quantum phase transition. Beyond $\delta>2$, superconductivity wins over the Kondo scale, and the impurity remains effectively unscreened.
In the next sections, we will use the Bethe Ansatz to study the low-energy physics of our model in each of its three boundary phases. We will show that all color excitations are gapped, while the $U(1)$ charge excitations remain gapless. As we vary the RG-invariant parameter $d$, the system moves between different boundary phases. We will examine the low-energy excitations in each phase. As shown in Ref.[44], the Hamiltonian (8) is integrable in the $J\to 0$ limit, and Ref.[34] demonstrates integrability in the $g\to 0$ limit. In Appendix A we construct the generic $N_{q}$‐particle eigenstates. We then show in Appendix B that the model remains integrable when both $g$ and $J$ are nonzero, and in Appendix C we derive the Bethe Ansatz equations. Moreover, the Bethe Ansatz equations are explicitly solved in Appendix D. Here, we shall summarize the results in all the passes in great detail.
@@ V.1 The Kondo Phase
[…]
@@ V.1.1 Fundamental bulk excitations in the Kondo phase
[…]
@@
[…]
@@
The most fundamental massive color excitation in the model is the bound state of a quark and an antiquark, obtained by introducing one hole into each of the rank-1 and rank-2 Bethe equations. Denote the hole rapidity in the rank-1 sector by $\theta$ and that in the rank-2 sector by $\vartheta$ (see Appendix D). Its energy is $E(\theta,\vartheta)=m\left(\cosh\frac{2\pi}{3}\theta+\cosh\frac{2\pi}{3}\vartheta\right)$, which when the rapidities are sent to infinity $\theta,\vartheta\to\infty$ attains its minimal value $E_{\min}=2m$. This excitation transforms in the Dynkin representation $(1,1)$, i.e. the 8-dimensional adjoint multiplet of $SU(3)$ with triality 0. The bare energy of the two-strings is exactly canceled by the backflow of the roots. The highest‐weight state has $Y_{\max}=1$ and $I_{3}=\tfrac{3}{2}$, and the eight states obtained by successive application of the lowering operators $E_{-\alpha_{1}}$ and $E_{-\alpha_{2}}$ give the usual meson‐octet diagram familiar from QCD, which we explicitly plot in the $Y$$I_{3}$ plane in Fig.3(a)a).
@@
When one adds a single two-string in each sector (i.e. $\lambda_{\alpha}\pm\frac{i}{2}$ in the rank-1 equation and $\mu_{k}\pm\frac{i}{2}$ in the rank-2 equation) on top of one hole per rank at rapidities $\theta$ (rank-1) and $\vartheta$ (rank-2), the resulting state is the meson singlet with Dynkin labels $(0,0)$(see Appendix D). Its energy is given by $E(\theta,\vartheta)=m\left(\cosh\frac{2\pi}{3}\theta+\cosh\frac{2\pi}{3}\vartheta\right)$, which attains its minimal value $E_{\min}=2m$ when rapidities $\theta,\vartheta\to\infty$. The energy contribution in the thermodynamic limits comes solely from the two holes since the bare energy of the two-strings is exactly canceled by the contribution from the backflow of the roots.
@@ V.2 The YSR Phase
A state where the impurity is screened by a boundary-localized bound mode can then be constructed by adding the boundary-string solution from the rank-1 Bethe Ansatz equations for the range $11/2<\delta<$ 1, and by adding both boundary-string solutions for the range $21<\delta<$ 2 (see Appendix D for details).
@@
such that the number of rank-1 roots is $M_{1}=\int\rho_{1}(\lambda)\mathrm{d}\lambda=\frac{2}{3}N_{q}$, and the number of rank-2 roots is $M_{2}=\int\rho_{2}(\mu)\mathrm{d}\mu=\frac{N_{q}}{3}$. Consequently, the state |$\ket{\psi_{1}}$ has Dynkin labels (1,0) and a $\mathbb{Z}_{3}$ triality of 1, which suggests that the impurity quark is unscreened.
Adding the purely imaginary root Eq. (34) in the regime $\frac{1}{2}<\delta<1$ and adding both imaginary roots Eq. (34) and Eq. (35) in the regime $21<\delta<$ 2, results in a state |$\ket{\psi_{\delta}}$ with Dynkin labels (0,0) with $\mathbb{Z}_{3}$ triality 0. In this state, the impurity quark is completely screened by the single particle localized bound mode. The change in root density of the rank-1 equation due to the presence of these roots is
@@
Figure 4: Energy of the boundary bound mode (i..e the energy difference between state |$\ket{\psi_{1}}$ and state |$\ket{\psi_{\delta}}$) as given by Eq. (4), showing that the mode’s energy is negative for $\frac{1}{2}<\delta<\frac{5}{4}$ and becomes positive for $\frac{5}{4}<\delta<1$.
Adding boundary strings to the state |$\ket{\psi_{1}}$ results in the ground state. When $41/2<\delta<5/$ 4, a single-particle bound mode screens the impurity. However, for $15/4<\delta<$ 1, this mid-gap state’s energy is higher than that of |$\ket{\psi_{1}}$. This indicates a boundary quantum phase transition in the model at $\delta=5/4$ due to the level crossing as shown in Fig. 4. For $\delta<5/4$, the impurity quark is screened in the ground state, whereas for $\delta>5/4$, it is unscreened.
All of the bulk excitations described in Sec. V.1.1 remain valid in this phase. In addition, there are boundary excitations involving the unscreened quark in the reference state |$\ket{\psi_{1}}$. For example, if one adds two holes at rapidities $\theta_{1}$ and $\theta_{2}$ into the rank-1 Bethe equation on top of |$\ket{\psi_{1}}$, one obtains the state |$\ket{\psi_{\theta_{1},\theta_{2}}}$. In this configuration, the unscreened quark, together with the two fundamental holes, carries Dynkin labels $(3,0)$ and triality 00. Its energy relative to |$\ket{\psi_{1}}$ is
$E_{\psi_{\theta_{1},\theta_{2}}}-E_{\ket{\psi_{1}}}=m\cosh\left(\frac{2\pi}{3}\theta_{1}\right)+m\cosh\left(\frac{2\pi}{3}\theta_{2}\right).$ (40)
@@
Similarly, creating a single hole at rapidity $\vartheta$ into the rank-2 Bethe equation on top of |$\ket{\psi_{1}}$ produces the state |$\ket{\psi_{\vartheta}}$. In this case, the excitation made up of an unscreened quark and the single hole carry Dynkin labels $(1,1)$ and triality 00. Their energy difference relative to |$\ket{\psi_{1}}$ is
$E_{\psi_{\vartheta}}-E_{\ket{\psi_{1}}}=m\cosh\left(\frac{2\pi}{3}\vartheta\right),$ (41)
@@
Thus far, all bulk excitations transform in the triality-00 (color‐singlet) sector, whereas at the boundary, the decoupling of the impurity quark permits triality‐$1$ states. We now demonstrate that, within the YSR phase and in certain parametric windows, even more exotic boundary excitations emerge. In the interval $1<\delta<\frac{3}{2}$, adding to the sea of real root solutions (of both the rank‐$1$ and rank‐$2$ Bethe equations) only the rank‐$1$ boundary string solutions, one obtains a boundary excitation carrying triality $2$. The total number of rank 1 roots including the one purely imaginary boundary string string solution is $M_{1}=1+\int\rho_{1}(\lambda)\mathrm{d}\lambda=\frac{1}{3}\left(2N_{b}+1\right)$ and the total number of rank 2 roots is $M_{2}=\int\rho_{2}(\mu)\mathrm{d}\mu=\frac{1}{3}\left(N_{b}-1\right)$ (see Appendix Sec.D.2). Thus, the Dynkin label for this state is $(p,q)=(0,1)$. The change in the density of rank 1 roots due to the boundary string solution is
@@
In the parametric regime $\delta>\frac{3}{2}$, adding only the rank 1 boundary string solution and a hole at position $\theta$ in the rank 1 solution leads to another triality 2 excitation with Dynkin label $(p,q)=(0,1)$ that transform in the antifundamental representation of $\rm SU(3)$. The change in the density of rank 1 roots due to the presence of the hole and the boundary string solution is
@@
In summary, in the intermediate YSR phase, realized for $\frac{1}{2}<\delta<2$, impurity screening is no longer driven by a many-body Kondo effect but instead occurs via the formation of a localized single-particle bound state near the impurity. The energy of this bound state,$|E_{\delta}|<m$, plays a decisive role in determining the ground state structure. For $\frac{1}{2}<\delta<\frac{5}{4}$, $E_{\delta}<0$ and hence occupying the bound state minimizes the energy, and the system favors a fully screened configuration corresponding to the SU(3) singlet state with Dynkin label $(p,q)=(0,0)$. Conversely, for the parametric regime, $\frac{5}{4}<\delta<2$, $E_{\delta}>0$ and hence the bound state remains unoccupied in the ground state, and the impurity quark exists as an unscreened degree of freedom, yielding a state with $(p,q)=(1,0)$. The transition between these two regimes occurs via a level crossing at $E_{\delta}=0$, which occurs at $\delta=\frac{5}{4}$, marking a first-order quantum phase transition. In addition to the bulk spectrum of color-singlet mesons and baryons, this phase also supports localized boundary excitations, wherein the unscreened impurity quark can bind with a bulk antiquark or two bulk quarks to form a meson-like or baryon-like state, respectively. Moreover, for $\delta>1$, one can realize an exotic triality-$2$ boundary excitation. In the interval $1<\delta<3/2,$ its energy lies strictly within the gap, $E_{\delta}<m$, whereas for $\delta>3/2$ it asymptotically coincides with the energy of a fundamental hole in the thermodynamic limit.
@@ VI The Unscreened Phase
In this regime, the two purely imaginary boundary‐string solutions from Eq. (34) and Eq. (35) remain valid. The ground state is |$\ket{\psi_{1}}$, in which the impurity is unscreened. Adding only the rank 1 boundary string solution leads to triality 2 excitation, discussed above with energy given by Eq.(45). If one adds both boundary strings to the Bethe‐root configuration of |$\ket{\psi_{1}}$, then the rank‐1 root density density are changed by
@@ VII Conclusion
In this work, we introduced a novel integrable quantum field theory, which is a toy model that captures the competition between color superconductivity and the color–Kondo effect in quark matter. In our model, the quarks are massless in the ultraviolet regime; however, an attractive $\mathrm{SU}(3)$ four‐fermion color density–density interaction between opposite‐chirality quarks drives the bulk coupling $g$ to strong coupling in the infrared, dynamically generating a mass gap $m$ (the color‐superconducting gap) via dynamical transmutation. Unlike QCD, our model contains no dynamical gluon bosons; instead, the attraction that gives rise to quasi-long-range superconducting order arises from a four‐fermion color–density–density interaction in bulk. In the presence of a heavy quark impurity the boundary coupling $J$ between the localized impurity and the bulk color density of light quarks renormalizes depending on the relative magnitudes of $J$ and $g$: for $J\gg g$, the system flows into a Kondo‐screened phase; for $J\sim g$, a Yu–Shiba–Rusinov bound state appears within the gap and screens the impurity; and for $g\gg J$, the impurity remains unscreened at low energy.
In the Kondo phase, we demonstrated that the impurity becomes completely screened below a dynamically generated Kondo scale $T_{K}$, despite the presence of a finite superconducting gap $m$. The low-energy spectrum is thus governed entirely by bulk color-singlet excitations like mesons and baryons with no boundary impurity excitations. In contrast to the Kondo phase, where the impurity is fully screened by many-body entanglement with the bulk, the intermediate YSR phase features screening via a localized single-particle bound state near the impurity. Consequently, the impurity can be unscreened by not ‘occupying’ the bound mode, which creates an excitation of energy $E_{\delta}<m$. Finally, there exists an unscreened phase in which the impurity can not be completely screened at any energy scale and, therefore, retains its full entropy of $\ln 3$ both in the ultraviolet and infrared limits.
@@ Appendix A $N_{q}$-particle eigenstate
In this section, we explain the process of determining the eigenstates of the Hamiltonian operator, denoted $\hat{H}$. Because the particle number operator $\hat{N}$ is a conserved quantity, satisfying the commutation relation $[\hat{H},\hat{N}]=0$, we can identify the eigenstates of $\hat{H}$ by exploring each subspace with a fixed number of bulk quarks $N_{q}$, independently. We initiate this analysis with the single-particle case, $N_{q}=1$, where the wave function can be represented as a superposition of plane waves as
|$\ket{k}=\sum_{a_{j},\sigma}\int_{-L}^{0}\mathrm{d}xe^{i\sigma kx}A^{\sigma}_{a_{1}a_{0}}\psi_{\sigma,a_{1}}^{\dagger}(x)\ket{0}\otimes\ket{a_{0}},$ (48)
where |$\ket{0}$ is the vacuum $\psi_{\sigma,a_{j}}\ket{0}=0$. Here $A^{\sigma}_{a_{j}a_{0}}$ are the amplitudes for the $j^{\mathrm{th}}$ itinerant fermion with chirality index $\sigma$ and color index $a_{j}$ scattering off the localized impurity carrying color $a_{0}$. Applying the Hamiltonian to the state |$\ket{k}$, Schrodinger’s equation fixes these amplitudes in the following form
@@
The identification of the impurity-electron S-matrix $S^{j0}$ completes the construction of the eigenstate |$\ket{k}$, which has eigenvalue $E=k$.
@@ A.1 Two particle sector
|$\ket{k_{i},k_{j}}=\sum_{\sigma,a}\int_{-L}^{0}\mathrm{d}^{2}xF_{a_{i}a_{j}a_{0}}^{\sigma_{i}\sigma_{j}}(x_{i},x_{j})e^{ik_{i}\sigma_{i}x_{i}+ik_{j}\sigma_{j}x_{j}}\psi_{\sigma_{i}a_{i}}^{\dagger}(x_{i})\psi_{\sigma_{j}a_{j}}^{\dagger}(x_{j})\ket{0}\otimes\ket{a_{0}},$ (53)
@@ Appendix B Integrability of the model
Consider a system of particles trapped between two reflective walls at $x=0$ and $x=-L$. We impose open boundary conditions to describe how particles behave at these walls. Let $\psi_{+}(x)$ be the wave function for particles moving right and $\psi_{-}(x)$ for those moving left. At the boundaries, we require:
@@
Now, imagine $N_{q}$ particles in this system, each with a quasimomentum $k_{j}$ (where $j=1,2,\ldots,N_{q}$). The wave function is a mix of plane waves, but the boundary conditions limit the possible $k_{j}$ values. Applying the condition at $x=-L$ to the $N_{q}$-particle wave function (as in Eq. (64)), we get:
@@
  • $e^{-2ik_{j}L}$ is the geometric phase a particle picks up traveling from $x=-L$ to $x=0$ and back—a round trip of distance $2L$.
@@
By leveraging the algebraic relations outlined in Eq. (65) through Eq. (67), which include the Yang-Baxter equation and associated consistency conditions, it can be rigorously established that the transfer matrices $Z_{j}$ and $Z_{k}$ commute for all $j,k$, satisfying the relation $[Z_{j},Z_{k}]=0$. This commutation property is a fundamental characteristic of integrable systems, indicating that all transfer matrices share a common set of eigenvectors and can thus be diagonalized simultaneously. This feature underpins the exact solvability of the model, enabling a precise determination of its spectral properties.

### v2

@@ IV Summary of Phases
We proceed to discuss the various phases the system possesses, as functions of the RG invariant parameter $d$. As shown in Fig. 4, for all values of $d\in\mathbb{R}$ and for $0<\delta<\frac{1}{2}$ the impurity quark is completely screened by a multiparticle cloud of bulk superconducting quarks. In this regime, the Kondo fluctuations overwhelm the bulk superconducting order, and in addition to the bulk mass gap $\Delta$, a new scale $T_{K}$, the Kondo temperature, is dynamically generated, which characterizes the screening of the impurity quark and the local impurity density of states, which attains the familiar Lorentzian form. As $\delta$ increases past $\frac{1}{2}$, superconducting pairing increases and competes with Kondo screening, causing a bound state to emerge below the bulk mass gap. In the range $\frac{1}{2}<\delta<\frac{5}{4}$, the bound‐state energy is negative, and the impurity is screened by a single‐particle mode. At $\delta=\frac{5}{4}$, the bound level crosses zero, and for $\frac{5}{4}<\delta<2$, its positive energy makes the unscreened impurity lower in energy, marking a first-order quantum phase transition. Beyond $\delta>2$, superconductivity wins over the Kondo scale, and the impurity remains effectively unscreened.
In the next sections, we will use the Bethe Ansatz to study the low-energy physics of our model in each of its three boundary phases. We will show that all color excitations are gapped, while the $U(1)$ charge excitations remain gapless. As we vary the RG-invariant parameter $d$, the system moves between different boundary phases. We will examine the low-energy excitations in each phase. As shown in Ref.[45], the Hamiltonian (8) is integrable in the $J\to 0$ limit, and Ref.[35] demonstrates integrability in the $g\to 0$ limit. In Appendix A we construct the generic $N_{q}$‐particle eigenstates. We then show in Appendix B that the model remains integrable when both $g$ and $J$ are nonzero, and in Appendix C we derive the Bethe Ansatz equations. Moreover, the Bethe Ansatz equations are explicitly solved in Appendix D. Here, we shall summarize the results in all the phases in great detail.
@@ V.1 The Kondo Phase
$\sum_{\beta=1}^{M_{1}}F(\lambda_{\beta})\to\int_{-\infty}^{\infty}F(\lambda^{\prime})\rho_{1}(\lambda^{\prime})d\lambda^{\prime},$
$\sum_{k=1}^{M_{2}}G(\mu_{k})\to\int_{-\infty}^{\infty}G(\mu^{\prime})\rho_{2}(\mu^{\prime})d\mu^{\prime}.$
@@ V.1.1 Fundamental bulk excitations in the Kondo phase
$I_{3}$YY(32,1)($\frac${3}{2},1)(12,1)($\frac${1}{2},1)(12,1)(-$\frac${1}{2},1)(32,1)(-$\frac${3}{2},1)(1,0)(1,0)(0,0)(0,0)(1,0)(-1,0)(12,1)($\frac${1}{2},-1)(12,1)(-$\frac${1}{2},-1)(0,2)(0,-2)
@@
$I_{3}$YY(0,2)(0,2)(12,1)($\frac${1}{2},1)(12,1)(-$\frac${1}{2},1)(1,0)(1,0)(0,0)(0,0)(1,0)(-1,0)(32,1)($\frac${3}{2},-$1)(\frac{1}{2},$-$1)(-\frac{1}{2},$-$1)(-\frac{3}{2},$-1)
@@
[…]
@@
When one adds a single two-string in each sector (i.e. $\lambda_{\alpha}\pm\frac{i}{2}$ in the rank-1 equation and $\mu_{k}\pm\frac{i}{2}$ in the rank-2 equation) on top of one hole per rank at rapidities $\theta$ (rank-1) and $\vartheta$ (rank-2), the resulting state is the meson singlet with Dynkin labels (0,0)(0,0)(see Appendix D). Its energy is given by $E(\theta,\vartheta)=m\left(\cosh\frac{2\pi}{3}\theta+\cosh\frac{2\pi}{3}\vartheta\right)$, which attains its minimal value $E_{\min}=2m$ when rapidities $\theta,\vartheta\to\infty$. The energy contribution in the thermodynamic limits comes solely from the two holes since the bare energy of the two-strings is exactly canceled by the contribution from the backflow of the roots.
@@ V.2 The YSR Phase
A state where the impurity is screened by a boundary-localized bound mode can then be constructed by adding the boundary-string solution from the rank-1 Bethe Ansatz equations for the range $1/2<\delta<1$, and by adding both boundary-string solutions for the range $1<\delta<2$ (see Appendix D for details).
@@
[…]
@@
[…]
@@
[…]
@@
Thus far, all bulk excitations transform in the triality-0 (color‐singlet) sector, whereas at the boundary, the decoupling of the impurity quark permits triality‐$1$ states. We now demonstrate that, within the YSR phase and in certain parametric windows, even more exotic boundary excitations emerge. In the interval $1<\delta<\frac{3}{2}$, adding to the sea of real root solutions (of both the rank‐$1$ and rank‐$2$ Bethe equations) only the rank‐$1$ boundary string solutions, one obtains a boundary excitation carrying triality $2$. The total number of rank 1 roots including the one purely imaginary boundary string solution is $M_{1}=1+\int\rho_{1}(\lambda)\mathrm{d}\lambda=\frac{1}{3}\left(2N_{b}+1\right)$ and the total number of rank 2 roots is $M_{2}=\int\rho_{2}(\mu)\mathrm{d}\mu=\frac{1}{3}\left(N_{b}-1\right)$ (see Appendix Sec.D.2). Thus, the Dynkin label for this state is $(p,q)=(0,1)$. The change in the density of rank 1 roots due to the boundary string solution is
@@
In the parametric regime $\delta>\frac{3}{2}$, adding only the rank 1 boundary string solution and a hole at position $\theta$ in the rank 1 solution leads to another triality 2 excitation with Dynkin label $(p,q)=(0,1)$ that transforms in the antifundamental representation of $\rm SU(3)$. The change in the density of rank 1 roots due to the presence of the hole and the boundary string solution is
@@
[…]
@@ VI The Unscreened Phase
[…]
@@ VII Conclusion
In this work, we introduced a novel integrable quantum field theory, which is a toy model that captures the competition between color-superconductivity and the color–Kondo effect in quark matter. In our model, the quarks are massless in the ultraviolet regime; however, an attractive $\mathrm{SU}(3)$ four‐fermion color density–density interaction between opposite‐chirality quarks drives the bulk coupling $g$ to strong coupling in the infrared, dynamically generating a mass gap $m$ (the color‐superconducting gap) via dynamical transmutation. Unlike QCD, our model contains no dynamical gluon bosons; instead, the attraction that gives rise to quasi-long-range superconducting order arises from a four‐fermion color–density–density interaction in bulk. In the presence of a heavy quark impurity the boundary coupling $J$ between the localized impurity and the bulk color density of light quarks renormalizes depending on the relative magnitudes of $J$ and $g$: for $J\gg g$, the system flows into a Kondo‐screened phase; for $J\sim g$, a Yu–Shiba–Rusinov bound state appears within the gap and screens the impurity; and for $g\gg J$, the impurity remains unscreened at low energy.
In the Kondo phase, we demonstrated that the impurity becomes completely screened below a dynamically generated Kondo scale $T_{K}$, despite the presence of a finite superconducting gap $m$. The low-energy spectrum is thus governed entirely by bulk color-singlet excitations like mesons and baryons with no boundary impurity excitations. In contrast to the Kondo phase, where the impurity is fully screened by many-body entanglement with the bulk, the intermediate YSR phase features screening via a localized single-particle bound state near the impurity. Consequently, the impurity can be unscreened by not ‘occupying’ the bound mode, which creates an excitation of energy $E_{\delta}<m$. Finally, there exists an unscreened phase in which the impurity cannot be completely screened at any energy scale and, therefore, retains its free moment entropy value of $\ln 3$, both in the ultraviolet and infrared limits.
@@ Appendix A $N_{q}$-particle eigenstate
In this section, we explain the process of determining the eigenstates of the Hamiltonian operator, denoted $\hat{H}$. Because the particle number operator $\hat{N}$ is a conserved quantity, satisfying the commutation relation $\hat{H},\hat{N}]=0$, we can identify the eigenstates of $\hat{H}$ by exploring each subspace with a fixed number of bulk quarks $N_{q}$, independently. We initiate this analysis with the single-particle case, $N_{q}=1$, where the wave function can be represented as a superposition of plane waves as
[…]
@@
[…]
@@ A.1 Two particle sector
[…]
@@ Appendix B Integrability of the model
Consider a system of particles trapped between two reflective walls at $x=0$ and x=Lx=-L. We impose open boundary conditions to describe how particles behave at these walls. Let $\psi_{+}(x)$ be the wave function for particles moving right and $\psi_{-}(x)$ for those moving left. At the boundaries, we require:
@@
Now, imagine $N_{q}$ particles in this system, each with a quasimomentum $k_{j}$ (where $j=1,2,\ldots,N_{q}$). The wave function is a mix of plane waves, but the boundary conditions limit the possible $k_{j}$ values. Applying the condition at x=Lx=-L to the $N_{q}$-particle wave function (as in Eq. (64)), we get:
@@
  • $e^{-2ik_{j}L}$ is the geometric phase a particle picks up traveling from x=Lx=-L to $x=0$ and back—a round trip of distance $2L$.
@@
By leveraging the algebraic relations outlined in Eq. (65) through Eq. (67), which include the Yang-Baxter equation and associated consistency conditions, it can be rigorously established that the transfer matrices $Z_{j}$ and $Z_{k}$ commute for all $j,k$, satisfying the relation $Z_{j},Z_{k}]=0$. This commutation property is a fundamental characteristic of integrable systems, indicating that all transfer matrices share a common set of eigenvectors and can thus be diagonalized simultaneously. This feature underpins the exact solvability of the model, enabling a precise determination of its spectral properties.